OpenStudy (anonymous):

whats the slope of the line that passes through the points (2, -3) and (5, 1)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \bf \frac{y_{1}-y_{2}}{x_{1}-x_{2}} }\) \(\huge\color{blue}{ \bf \frac{(-3)-(1)}{(2)-(5)} }\) \(\huge\color{blue}{ \bf \frac{-4 ~}{-3~} }\) \(\huge\color{blue}{ \bf \frac{~4 ~}{~3~} }\)